Consider the reaction : 2MnO4- + 5H2O2 + 6H+ --> 2Mn2+ +5O2 +8H2O. If it takes 36.44 mL of 0.01652 M KMnO4 solution are required to oxidize 25.00 mL of a H2O2 solution, Calculate the molarity of the H2O2 solution.State your final answer using 4 Sigfigs.
Answer #1

from given equation,
2 moles of KMnO4 reacts with 5 moles of H2O2

so,
moles of H2O2 required = (5/2)* moles of KMnO4
M(H2O2)*V(H2O2) = (5/2)*M(KMnO4)*V(KMnO4)
M(H2O2) * 25.00 = (5/2)*0.01652 * 36.44
M(H2O2) = 0.06020 M
Answer: 0.06020 M